# Multimodal Deep Learning for Stroke Prediction and Detection using Retinal Imaging and Clinical Data

### v1

@@ Abstract
Stroke is a major public health problem, affecting millions worldwide. Deep learning has recently demonstrated promise for enhancing the diagnosis and risk prediction of stroke. However, existing methods rely on costly medical imaging modalities, such as computed tomography. Recent studies suggest that retinal imaging could offer a cost-effective alternative for cerebrovascular health assessment due to the shared clinical pathways between the retina and the brain. Hence, this study explores the impact of leveraging retinal images and clinical data for stroke detection and risk prediction. We propose a multimodal deep neural network that processes Optical Coherence Tomography (OCT) and infrared reflectance retinal scans, combined with clinical data, such as demographics, vital signs, and diagnosis codes. We pretrained our model using a self-supervised learning framework using a real-world dataset consisting of $37$ k scans, and then fine-tuned and evaluated the model using a smaller labeled subset. Our empirical findings establish the predictive ability of the considered modalities in detecting lasting effects in the retina associated with acute stroke and forecasting future risk within a specific time horizon. The experimental results demonstrate the effectiveness of our proposed framework by achieving $5$% AUROC improvement as compared to the unimodal image-only baseline, and $8$% improvement compared to an existing state-of-the-art foundation model. In conclusion, our study highlights the potential of retinal imaging in identifying high-risk patients and improving long-term outcomes.
@@
©Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works. This file corresponds to the accepted version of the manuscript published in 47th Annual International Conference of the IEEE Engineering in Medicine and Biology Society. DOI: ¡DOI No.¿
@@ I INTRODUCTION
Stroke is one of the leading causes of death and long-term disability worldwide, especially among the elderly [1]. The burden of stroke is a pressing global public health problem, with a remarkable increase observed between 1990 and 2019 [2], with low- and middle-income countries accounting for the largest proportion [3]. This underscores the critical need for robust stroke risk prediction models that enable early prevention among high-risk individuals.
Stroke occurs due to cerebral artery occlusion, also known as ischemic stroke, representing around $80\%$ of cases, or cerebral artery rupture, also known as hemorrhagic stroke [4]. Common stroke risk factors include history of other disease, such as diabetes, atrial fibrillation, hypertension, and renal failure, or lifestyle factors, such as tobacco use [5]. Stroke is diagnosed based on the sudden onset of associated symptoms, followed by rapid deterioration of the patient [6]. This then calls for swift interventions to reduce brain damage and mitigate effects [7].
@@
Computed Tomography (CT) and Magnetic Resonance Imaging (MRI) are considered the gold standards for stroke diagnosis following the onset of symptoms [8]. Both techniques are costly, relatively time-consuming, and subject to availability in healthcare centers. This signifies the need for exploring other effective techniques suitable for stroke risk assessment, which are lower in cost and can be quickly acquired in advance of the symptoms onset.
@@
The retina and the brain are known to share anatomical, physiological, and embryological origins [9]. This renders the retina as a potential candidate for understanding changes occurring in the brain [10]. Furthermore, recent work in neuro-ophthalmology has illustrated that Optical Coherence Tomography (OCT) can encode important information that reflects neural changes in the brain [11]. OCT is a fast, noninvasive, and cost-efficient technique that provides high-resolution in-vivo images of retinal tissue [12]. Recent studies validated the relevance of OCT imaging in diagnosing various neurological disease [13], such as Alzheimer’s. This work falls under the umbrella of an emerging area of research called oculomics, which aims to improve our understanding of systemic health and disease through the use of high-resolution retinal imaging and data science techniques [14].
@@
Artificial Intelligence (AI), particularly machine learning and deep learning, has significantly advanced healthcare to improve patient diagnosis, prognosis, and treatment. In ophthalmology, the increasing complexity of multimodal datasets and AI techniques offer a strong foundation for exploring the eye-body relationship. For example, several studies focused on the coupled use of retinal images and machine learning for cardiovascular [15] and neuro-degenerative disease [16]. However, existing studies that focus on the analysis of retinal images for stroke using machine learning are relatively limited. Existing work focuses on deriving retinal biomarkers for statistical machine learning models, such as regression and tree-based models, from color fundus photography [17, 18, 19, 20] and OCT angiography [21].
@@
A few studies attempted to develop deep learning models for stroke prediction using retinal images. For example, [22] trained three different Convolutional Neural Networks (CNN) for one-year stroke risk prediction using fundus photos of different wave lengths. Another work [23] developed a CNN using OCT angiography scans to distinguish between retinal images from stroke and non-stroke patients. Lastly, one study [24] proposed a foundation model (RetFound) for various systemic disease prediction, including ischemic stroke, and trained the model using a large dataset of color fundus photos and OCT separately. The model showed promising performance in predicting three-year stroke risk on the internal validation set but not on the external test set.
@@ II-D Prediction Model
As we had a relatively small labeled dataset to train RetStroke, we adopted a two-stage training strategy: self-supervised pre-training of the visual encoder with unlabeled data followed by downstream fine-tuning with the labeled dataset. For self-supervised pre-training, we adopted SimCLR [25], which is an unsupervised pre-training framework that uses the contrastive loss to learn high quality representations. The contrastive loss allows the model to learn representations by maximizing the similarity between positive pairs and minimizing it with negative pairs:
@@
For contrastive model pre-training, we developed two OCT datasets and one infrared reflectance dataset. Since the OCT scan is of volumetric nature, where a single eye scan contains multiple slices ranging in $(25-49)$ scans, we treated the slices as independent samples, which resulted in a dataset containing $1.1$ million image. The second pre-training OCT dataset consisted of the mid-slice only of each OCT volume, while the infrared reflectance dataset consisted of a single image from each study in the dataset. We split the pre-training datasets based on ratio of $90/10$ training/validation splits. After pretraining, we load the pre-trained weights to initialize the visual encoder encoder, $f_{oct}$, of RetStroke and we fine-tuned the entire architecture using the stroke labels.
@@ II-E Experimental Setup
To setup SimCLR for pretraining, we followed the same augmentations used in the original paper, including random resized crop, color jittering, Gaussian blur, and random vertical and horizontal flip (referred to as harsh augmentations). We also computed dataset specific statistics and use them for data normalization. We used AdamW [26] as an optimizer and a batch size of $256$. We also used cosine annealing for the learning rate scheduler. For $\tau$, we set the initial value as $0.5$ and fix it as $0.1$ if its value goes below this threshold during training. We pre-train each model for $200$ epochs with early stopping, using a patience of $10$ epochs and a minimum change of $61e-$ 6 in the validation loss. We also optimize the learning rate and weight decay with values in the intervals $[1e-6,1e-5]$ and $[0,1e-1]$, respectively, for the pre-training experiments using random search with 10 runs for each experiment.
@@
To optimize the hyperparameters of RetStroke, we performed Bayesian hyperparameter search with K-fold cross-validation. We used the Area Under the Receiver Operating Characteristic curve (AUROC) to select the best models using the validation set. We set the number of folds to five in all experiments. The optimized hyperparameters include the learning rate with values in $[1e-6,5e-5]$, weight decay with values in $[0,1e-3]$, batch size with values in $[64,128,256]$, and augmentations with values in [simple, harsh]. We also set the number of epochs during hyperparameter tuning to $100$ with a patience of $10$ epochs. We performed $100$ optimization trials for each experiment. After hyperparameter tuning, we selected the model with the best cross-validation AUROC and evaluated on the test set. We used Nvidia A100 GPUs for all experiments.
@@ III RESULTS
To better understand the performance of our proposed model, RetStroke, we conduct a subgroup analysis considering patient age, comorbidities, and stroke subtype. As shown in Figure 5, RetStroke achieves the best performance in the $(40-60)$ years subgroup, with an AUROC score of $0.740$ and $0.690$ for OCT and infrared, respectively. For the younger age group $(<40)$, the performance of RetStroke slightly degrades, achieving AUROC score of $0.700$ and $0.670$ for OCT and infrared, respectively. Finally, RetStroke shows a relatively low performance with respect to the elderly age group $(>60)$ compared to the other age groups, around $0.640$. We note that the mean age of the patient cohort is 57.8.
Figure 4 depicts RetStroke performance with respect to stroke subtypes. We observe that RetStroke performs better with respect to ischemic subtypes, including IS and TIA, with an AUROC score around $0.710$ for both subtypes using OCT and $0.650$ using infrared. The model achieves lower AUROC scores for ICH ($0.61-0.63$). This could be justified by the high prevalence of ischemic stroke ($80\%$) compared to the hemorrhagic stroke ($20\%$).
@@
Figure 3 illustrates RetStroke’s performance concerning patients’ historical comorbidities prior to scan acquisition. The best performance is observed among patients with blood-related diseases (D50-D89) using OCT scans, achieving an AUROC score of 0.819, while performance is significantly lower with infrared scans (0.711) for the same comorbidity. For nervous system (G00-G99) and circulatory system (I00-I99) disease, RetStroke achieves similar performance, with AUROC scores around 0.760 using OCT scans history of stroke-related diseases. Conversely, RetStroke exhibits the lowest performance among patients with eye disease (H00-H59), with AUROC scores of 0.686 using OCT scans and 0.640 using infrared scans. This finding aligns with previous research indicating that the presence of ophthalmic diseases degrades the model’s predictive performance [10, 27].
@@
We also investigated the performance of RetStroke with respect to stroke prediction and detection of lasting events, as shown in Table III. With infrared scans, RetStroke achieves a higher performance in stroke risk prediction especially in the first six months, with AUROC of $0.774$ and $0.766$ for the time periods $<90$ and $<180$, respectively. As the time period increases, performance remarkably decreases, with AUROC reaching below $0.710$ for $<270$ and $<365$ days. On the other hand, the ability of RetStroke to detect lasting effects of stroke is relatively lower with infrared compared to risk prediction, with an AUROC ranging between $0.620$ and $0.640$.
With the OCT scans, RetStroke achieves an AUROC of $0.723$ and $0.736$ for risk prediction within $<90$ and $<180$ days, respectively. The performance similarly decreases within the longer time horizons with AUROC scores dropping below $0.71$. For the detection of lasting effects, RetStroke has a similar performance across the first three time horizons with AUROC $0.705-0.719$.. One interesting observation is that RetStroke generally performs better in the risk prediction task using the infrared modality. While for lasting effect, RetStroke performs significantly better with OCT than infrared across all time horizons.
@@ IV DISCUSSION & CONCLUSIONS
In this paper, we present RetStroke, a clinically informed framework that is able to leverage retinal images for stroke detection and prediction. The main advantage of RetStroke is that it is multimodal by nature, since it incorporates patient’s comorbidities, vital-sign measurements, and demographics during training and inference to enhance its accuracy. RetStroke demonstrated a superior performance gain as compared to unimodal models trained on imaging data only, such as OCT and infrared. Furthermore, Retstroke achieve a better performance than an existing foundation model, RetFound [24], which consists of around $300$ million parameters and was pre-trained using a significantly larger OCT dataset in the unimodal setting. We investigated incorporating the clinical information in the existing foundation model RetFound and showed that such an adaptation leads to significant improvement in performance. These results highlight the value of the clinical information utilized in our proposed framework. Additionally, our evaluation of RetStroke across diverse patients subgroups including age groups, stroke subtypes and comorbidity indicates its robustness. Overall, RetStroke can be regarded as a novel and unique addition to the field of oculomics to enhance our understanding of an important disease such as stroke.
@@
On the other hand, this study has some limitations that point to promising research directions. First, although our method is compared against external baselines, RetStroke has only been evaluated on an internal private dataset, raising concerns about its generalizability. However, accessing such data is challenging due to healthcare data privacy and regulatory constraints, necessitating further efforts to test the model on external datasets as well as additional baselines. Second, the dataset used for RetStroke development is relatively small (183 patients), which limits the model’s ability to learn stroke-specific features, thereby affecting performance. Addressing this issue requires further research. Possible research directions on this point include training with loss functions that handle class imbalance such as focal loss [28], or applying rare-class oversampling strategies. Third, we define our risk prediction time-frame as a maximum of one year, but longer periods could reveal different insights. Lastly, although RetStroke was developed specifically for stroke, applying it to a broader range of diseases, such as cardiovascular and neurodegenerative conditions, could enhance its clinical utility. Investigating RetStroke in these areas is a pressing need. Overcoming these limitations will contribute to advancing this important field of research.

### v2

@@ Abstract
Stroke is a major public health problem, affecting millions worldwide. Deep learning has recently demonstrated promise for enhancing the diagnosis and risk prediction of stroke. However, existing methods rely on costly medical imaging modalities, such as computed tomography. Recent studies suggest that retinal imaging could offer a cost-effective alternative for cerebrovascular health assessment due to the shared clinical pathways between the retina and the brain. Hence, this study explores the impact of leveraging retinal images and clinical data for stroke detection and risk prediction. We propose a multimodal deep neural network that processes Optical Coherence Tomography (OCT) and infrared reflectance retinal scans, combined with clinical data, such as demographics, vital signs, and diagnosis codes. We pretrained our model using a self-supervised learning framework using a real-world dataset consisting of $37$ k scans, and then fine-tuned and evaluated the model using a smaller labeled subset. Our empirical findings establish the predictive ability of the considered modalities in detecting lasting effects in the retina associated with acute stroke and forecasting future risk within a specific time horizon. The experimental results demonstrate the effectiveness of our proposed framework by achieving 55% AUROC improvement as compared to the unimodal image-only baseline, and 88% improvement compared to an existing state-of-the-art foundation model. In conclusion, our study highlights the potential of retinal imaging in identifying high-risk patients and improving long-term outcomes.
@@
©Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works. This file corresponds to the accepted version of the manuscript published in 47th Annual International Conference of the IEEE Engineering in Medicine and Biology Society. DOI: 10.1109
@@ I INTRODUCTION
Stroke is one of the leading causes of death and long-term disability worldwide, especially among the elderly [feigin2021global]. The burden of stroke is a pressing global public health problem, with a remarkable increase observed between 1990 and 2019 [feigin2022world], with low- and middle-income countries accounting for the largest proportion [katan2018global]. This underscores the critical need for robust stroke risk prediction models that enable early prevention among high-risk individuals.
Stroke occurs due to cerebral artery occlusion, also known as ischemic stroke, representing around $80\%$ of cases, or cerebral artery rupture, also known as hemorrhagic stroke [tobin2014neurogenesis]. Common stroke risk factors include history of other disease, such as diabetes, atrial fibrillation, hypertension, and renal failure, or lifestyle factors, such as tobacco use [boehme2017stroke]. Stroke is diagnosed based on the sudden onset of associated symptoms, followed by rapid deterioration of the patient [mosley2007stroke]. This then calls for swift interventions to reduce brain damage and mitigate effects [von2019time].
@@
Computed Tomography (CT) and Magnetic Resonance Imaging (MRI) are considered the gold standards for stroke diagnosis following the onset of symptoms [merino2010imaging]. Both techniques are costly, relatively time-consuming, and subject to availability in healthcare centers. This signifies the need for exploring other effective techniques suitable for stroke risk assessment, which are lower in cost and can be quickly acquired in advance of the symptoms onset.
@@
The retina and the brain are known to share anatomical, physiological, and embryological origins [london2013retina]. This renders the retina as a potential candidate for understanding changes occurring in the brain [girach2024retinal]. Furthermore, recent work in neuro-ophthalmology has illustrated that Optical Coherence Tomography (OCT) can encode important information that reflects neural changes in the brain [lamirel2009use]. OCT is a fast, noninvasive, and cost-efficient technique that provides high-resolution in-vivo images of retinal tissue [xie2022use]. Recent studies validated the relevance of OCT imaging in diagnosing various neurological disease [awais2017classification], such as Alzheimer’s. This work falls under the umbrella of an emerging area of research called oculomics, which aims to improve our understanding of systemic health and disease through the use of high-resolution retinal imaging and data science techniques [wagner2020insights].
@@
Artificial Intelligence (AI), particularly machine learning and deep learning, has significantly advanced healthcare to improve patient diagnosis, prognosis, and treatment. In ophthalmology, the increasing complexity of multimodal datasets and AI techniques offer a strong foundation for exploring the eye-body relationship. For example, several studies focused on the coupled use of retinal images and machine learning for cardiovascular [hu2023systematic] and neuro-degenerative disease [bahr2024deep]. However, existing studies that focus on the analysis of retinal images for stroke using machine learning are relatively limited. Existing work focuses on deriving retinal biomarkers for statistical machine learning models, such as regression and tree-based models, from color fundus photography [messica2024enhancing, yuanyuan2020comparison, zhu2022retinal, rudnicka2022artificial] and OCT angiography [liang2022retinal].
@@
A few studies attempted to develop deep learning models for stroke prediction using retinal images. For example, [li2023predicting] trained three different Convolutional Neural Networks (CNN) for one-year stroke risk prediction using fundus photos of different wave lengths. Another work [xiong2024association] developed a CNN using OCT angiography scans to distinguish between retinal images from stroke and non-stroke patients. Lastly, one study [zhou2023foundation] proposed a foundation model (RetFound) for various systemic disease prediction, including ischemic stroke, and trained the model using a large dataset of color fundus photos and OCT separately. The model showed promising performance in predicting three-year stroke risk on the internal validation set but not on the external test set.
@@ II-D Prediction Model
As we had a relatively small labeled dataset to train RetStroke, we adopted a two-stage training strategy: self-supervised pre-training of the visual encoder with unlabeled data followed by downstream fine-tuning with the labeled dataset. For self-supervised pre-training, we adopted SimCLR [chen2020simple], which is an unsupervised pre-training framework that uses the contrastive loss to learn high quality representations. The contrastive loss allows the model to learn representations by maximizing the similarity between positive pairs and minimizing it with negative pairs:
@@
For contrastive model pre-training, we developed two OCT datasets and one infrared reflectance dataset. Since the OCT scan is of volumetric nature, where a single eye scan contains multiple slices ranging in (2549)(25-49) scans, we treated the slices as independent samples, which resulted in a dataset containing $1.1$ million image. The second pre-training OCT dataset consisted of the mid-slice only of each OCT volume, while the infrared reflectance dataset consisted of a single image from each study in the dataset. We split the pre-training datasets based on ratio of $90/10$ training/validation splits. After pretraining, we load the pre-trained weights to initialize the visual encoder encoder, $f_{oct}$, of RetStroke and we fine-tuned the entire architecture using the stroke labels.
@@ II-E Experimental Setup
To setup SimCLR for pretraining, we followed the same augmentations used in the original paper, including random resized crop, color jittering, Gaussian blur, and random vertical and horizontal flip (referred to as harsh augmentations). We also computed dataset specific statistics and use them for data normalization. We used AdamW [loshchilov2017decoupled] as an optimizer and a batch size of $256$. We also used cosine annealing for the learning rate scheduler. For $\tau$, we set the initial value as $0.5$ and fix it as $0.1$ if its value goes below this threshold during training. We pre-train each model for $200$ epochs with early stopping, using a patience of $10$ epochs and a minimum change of 1e61e-6 in the validation loss. We also optimize the learning rate and weight decay with values in the intervals [1e6,1e5][1e-6,1e-5] and [0,1e1][0,1e-1], respectively, for the pre-training experiments using random search with 10 runs for each experiment.
@@
To optimize the hyperparameters of RetStroke, we performed Bayesian hyperparameter search with K-fold cross-validation. We used the Area Under the Receiver Operating Characteristic curve (AUROC) to select the best models using the validation set. We set the number of folds to five in all experiments. The optimized hyperparameters include the learning rate with values in [1e6,5e5][1e-6,5e-5], weight decay with values in [0,1e3][0,1e-3], batch size with values in $[64,128,256]$, and augmentations with values in [simple, harsh]. We also set the number of epochs during hyperparameter tuning to $100$ with a patience of $10$ epochs. We performed $100$ optimization trials for each experiment. After hyperparameter tuning, we selected the model with the best cross-validation AUROC and evaluated on the test set. We used Nvidia A100 GPUs for all experiments.
@@ III RESULTS
To better understand the performance of our proposed model, RetStroke, we conduct a subgroup analysis considering patient age, comorbidities, and stroke subtype. As shown in Figure 5, RetStroke achieves the best performance in the (4060)(40-60) years subgroup, with an AUROC score of $0.740$ and $0.690$ for OCT and infrared, respectively. For the younger age group $(<40)$, the performance of RetStroke slightly degrades, achieving AUROC score of $0.700$ and $0.670$ for OCT and infrared, respectively. Finally, RetStroke shows a relatively low performance with respect to the elderly age group $(>60)$ compared to the other age groups, around $0.640$. We note that the mean age of the patient cohort is 57.8.
Figure 4 depicts RetStroke performance with respect to stroke subtypes. We observe that RetStroke performs better with respect to ischemic subtypes, including IS and TIA, with an AUROC score around $0.710$ for both subtypes using OCT and $0.650$ using infrared. The model achieves lower AUROC scores for ICH (0.610.630.61-0.63). This could be justified by the high prevalence of ischemic stroke ($80\%$) compared to the hemorrhagic stroke ($20\%$).
@@
Figure 3 illustrates RetStroke’s performance concerning patients’ historical comorbidities prior to scan acquisition. The best performance is observed among patients with blood-related diseases (D50-D89) using OCT scans, achieving an AUROC score of 0.819, while performance is significantly lower with infrared scans (0.711) for the same comorbidity. For nervous system (G00-G99) and circulatory system (I00-I99) disease, RetStroke achieves similar performance, with AUROC scores around 0.760 using OCT scans history of stroke-related diseases. Conversely, RetStroke exhibits the lowest performance among patients with eye disease (H00-H59), with AUROC scores of 0.686 using OCT scans and 0.640 using infrared scans. This finding aligns with previous research indicating that the presence of ophthalmic diseases degrades the model’s predictive performance [girach2024retinal, diaz2022predicting].
@@
We also investigated the performance of RetStroke with respect to stroke prediction and detection of lasting events, as shown in Table III. With infrared scans, RetStroke achieves a higher performance in stroke risk prediction especially in the first six months, with AUROC of $0.774$ and $0.766$ for the time periods <90<90 and <180<180, respectively. As the time period increases, performance remarkably decreases, with AUROC reaching below $0.710$ for <270<270 and <365<365 days. On the other hand, the ability of RetStroke to detect lasting effects of stroke is relatively lower with infrared compared to risk prediction, with an AUROC ranging between $0.620$ and $0.640$.
With the OCT scans, RetStroke achieves an AUROC of $0.723$ and $0.736$ for risk prediction within <90<90 and <180<180 days, respectively. The performance similarly decreases within the longer time horizons with AUROC scores dropping below $0.71$. For the detection of lasting effects, RetStroke has a similar performance across the first three time horizons with AUROC 0.7050.7190.705-0.719.. One interesting observation is that RetStroke generally performs better in the risk prediction task using the infrared modality. While for lasting effect, RetStroke performs significantly better with OCT than infrared across all time horizons.
@@ IV DISCUSSION & CONCLUSIONS
In this paper, we present RetStroke, a clinically informed framework that is able to leverage retinal images for stroke detection and prediction. The main advantage of RetStroke is that it is multimodal by nature, since it incorporates patient’s comorbidities, vital-sign measurements, and demographics during training and inference to enhance its accuracy. RetStroke demonstrated a superior performance gain as compared to unimodal models trained on imaging data only, such as OCT and infrared. Furthermore, Retstroke achieve a better performance than an existing foundation model, RetFound [zhou2023foundation], which consists of around $300$ million parameters and was pre-trained using a significantly larger OCT dataset in the unimodal setting. We investigated incorporating the clinical information in the existing foundation model RetFound and showed that such an adaptation leads to significant improvement in performance. These results highlight the value of the clinical information utilized in our proposed framework. Additionally, our evaluation of RetStroke across diverse patients subgroups including age groups, stroke subtypes and comorbidity indicates its robustness. Overall, RetStroke can be regarded as a novel and unique addition to the field of oculomics to enhance our understanding of an important disease such as stroke.
@@
On the other hand, this study has some limitations that point to promising research directions. First, although our method is compared against external baselines, RetStroke has only been evaluated on an internal private dataset, raising concerns about its generalizability. However, accessing such data is challenging due to healthcare data privacy and regulatory constraints, necessitating further efforts to test the model on external datasets as well as additional baselines. Second, the dataset used for RetStroke development is relatively small (183 patients), which limits the model’s ability to learn stroke-specific features, thereby affecting performance. Addressing this issue requires further research. Possible research directions on this point include training with loss functions that handle class imbalance such as focal loss [lin2017focal], or applying rare-class oversampling strategies. Third, we define our risk prediction time-frame as a maximum of one year, but longer periods could reveal different insights. Lastly, although RetStroke was developed specifically for stroke, applying it to a broader range of diseases, such as cardiovascular and neurodegenerative conditions, could enhance its clinical utility. Investigating RetStroke in these areas is a pressing need. Overcoming these limitations will contribute to advancing this important field of research.